# Revisiting Synthetic Human Trajectories: Imitative Generation and Benchmarks Beyond Datasaurus

### v1

@@ 5.2.2. Location Recommendation Task
(LocRec) suggests new (previously unvisited) locations for users by modeling user preferences on locations from trajectory data (Ye et al., 2010). Following the default setting of (Zhao et al., 2021), we transform a trajectory dataset into a set of (user, location, visit_count) triplets and then split them into training/valid/test datasets under a ratio of 8:1:1. To discount the impact of the specific techniques and metrics, we consider five popular recommendation algorithms, i.e., BPR (Rendle et al., 2012), DMF (Xue et al., 2017), LightGCN (He et al., 2020), MultiVAE (Liang et al., 2018), and NeuMF (He et al., 2017) (details in Appendix E), and report their performance on Mean Reciprocal Rank@N (MRR@N), Normalized Discounted Cumulative Gain@N (NDCG@N), hit@N (where N = 5 and 10). Finally, we compare the paired performance discrepancy between the real and generated trajectories (i.e., 30 paired results from five algorithms and six metrics each) using MAPE and MSPE.
@@ 5.2.3. Next Location Prediction Task
(NexLoc) forecasts a user’s next location in the future by learning the sequential mobility patterns from historical user trajectories (Noulas et al., 2012). Specifically, for one dataset, we chronologically split each trajectory into training/valid/test trajectories under a ratio of 8:1:1. We also consider five sequence prediction algorithms, i.e., FPMC (Rendle et al., 2010), BERT4Rec (Sun et al., 2019), Caser (Tang and Wang, 2018), SRGNN (Wu et al., 2019), and SASRec (Kang and McAuley, 2018) (details in Appendix F), and report their performance on MRR@N, NDCG@N, hit@N (where N = 5 and 10). Finally, we compare the paired performance discrepancy between the real and generated trajectories using MAPE and MSPE.
@@ 5.2.5. Epidemic Simulation Task
(EpiSim) simulates the epidemic spreading over a contact network characterizing the spatiotemporal contact patterns of user trajectories (Verma et al., 2021). The contact network is extracted from a trajectory dataset as a dynamic graph of users, and a directed edge from a user $p$ to a user $q$ represents that $p$’s visit precedes $q$’s visit to the same location during a day, indicating a potential chance of epidemic spreading from $p$ to $q$. Following the setting of recent works (Feng et al., 2020; Yuan et al., 2022; Lai et al., 2020), we adopt the Susceptible–Exposed–Infected–Recovered (SEIR) model. We consider the simulation of COVID-19 using the parameters suggested by (Feng et al., 2020; Yuan et al., 2022) and influenza using the parameters suggested by (Brauer et al., 2019) (details in Appendix G). We randomly select 50 individuals as exposed individuals, simulate the spread of the epidemic, and report the daily counts of Exposed, Infectious, and Recovered individuals as three metrics. We report the average results of 10 repeated simulations to discount the impact of the random selection of initially exposed individuals. Finally, we compare the performance discrepancy between the real and generated trajectories using MAPE and MSPE.
@@ 5.3. Statistical & Distributional Similarity
Table 2 shows the performance comparison of MIRAGE and baselines on the three datasets. We observe that MIRAGE consistently achieves the best performance with the lowest JSD, yielding 59.0%, 68.9%, and 71.5% improvement (reduction on JSD) over the best-performing baselines on TKY, IST, and NYC datasets, respectively. However, even though we selected four similarity metrics covering three different aspects of spatial (Distance and Radius), temporal (Interval), and user (DailyLoc) distributions, these metrics cannot fully reflect the ultimate utility of generated trajectories in supporting downstream tasks. For example, the consistent superiority of MIRAGE over baselines on these similarity metrics is still biased, because MIRAGE indeed underperforms some baselines in a few cases in our task-based evaluation, as we discuss below.
@@ 5.4. Task-Based Evaluation Performance
Table 3 shows the performance in our proposed task-based evaluation on MAPE (similar results on MSPE, shown in Appendix H). Each entry in this Table is the MAPE averaging over all metrics of all algorithms solving a task on a dataset, to average out the biases of individual techniques and metrics. For example, the MAPE of MIRAGE in the LocRec task on the TKY dataset is 0.3485, which is computed from 30 paired (real and generated) results from five algorithms and six metrics used in the LocRec task.
@@ 5.5. Ablation Study
We conduct an ablation study on our proposed method MIRAGE, considering the following two variants. MIRAGE-noTPP is a variant of MIRAGE without the neural TPP, where we define a regression task to predict the next inter-event from the time context vector $c_{i}^{\tau}$. MIRAGE-noEPR is a variant of MIRAGE without neural EPR, where we directly sample a location from a categorical distribution conditioned on the location context vector $c_{i}^{l}$. Tables 4 and 5 show the results on similarities and task-based evaluation on MAPE, respectively (similar results on MSPE in Appendix H).
We observe that MIRAGE consistently outperforms MIRAGE-noTPP on both similarities and task-based evaluation by 63.8% and 31.2% (on average over tasks and datasets), respectively, showing the effectiveness of neural TPPs modeling the event stochasticity of human trajectories. Second, MIRAGE outperforms MIRAGE-noEPR in most cases, with an improvement of 50.3% and 19.2% on similarities and task-based evaluation, respectively, which verifies the usefulness of our neural EPR model. In addition, to further show the utility of the EPR model, we plot the empirical returning probability of users over time, which is defined as the probability of a user returning to a location a certain period (temporal distance) after the user’s first presence at the location (Gonzalez et al., 2008; Yang et al., 2020a). Figure 2 shows the plots of both real and generated data on NYC. We see that the real trajectory exhibits strong periodicity, which can be well imitated by MIRAGE but not by MIRAGE-noEPR. The plots for baselines are also available in Appendix I. The superiority of MIRAGE is also validated on the return probability over sequence lengths in Appendix J.
@@ 6. Conclusions and Future Works
In this paper, by revisiting the existing human trajectory generative models, we identify their limitations in focusing on the summary statistics and distributional similarities between real and generated trajectories, which could lead to intrinsic biases in both generative model design and benchmarks of the generated trajectories. Against this background, we propose MIRAGE, a huMan-Imitative tRAjectory GenErative model designed as an intensity-free neural Temporal Point Process integrating a neural Exploration and Preferential Return model to imitate the human decision-making process in trajectory generation. Meanwhile, we also propose a comprehensive task-based evaluation protocol to systematically benchmark trajectory generative models on four typical downstream tasks, integrating multiple techniques and evaluation metrics for each task, to assess the ultimate utility of the generated trajectories. The evaluation results show that MIRAGE-generated trajectory data not only achieves the best statistical and distributional similarities with 59.0-71.5% improvement but also yields the best performance in the task-based evaluation with 10.9-33.4% improvement.
@@ Appendix A Biases in downstream tasks
We show the performance variation of different methods solving the same downstream task (even average over multiple metrics) in our task-based evaluation on MAPE in Table 6. Taking MoveSim and VOLUNTEER as examples, MoveSim is better than VOLUNTEER when benchmarking using FPMC or SASRec in this task, while we have the opposite results if benchmarking using BERT4Rec, Caser, or SRGNN.
Appendix B The Visualization of the Return Probability over Different Time Distances
Figure 3 shows the learnt return probability $p^{*}(l_{i+1}|\tau_{i+1},return)$ over different time distances (hour granularities in a week) from each dataset. We see a clear daily return pattern, indicating that our model can effectively capture the periodicity encoded in human trajectories.
@@ Appendix B The Visualization of the Return Probability over Different Time Distances
Appendix C The Complexity of MIRAGE
@@
The complexity of our MIRAGE is discussed as follows. Let $n$ denote the sequence length, $d$ is the GRU hidden state size, and $z$ is the latent size for user VAE. Our time decoder with intensity-free TPP and category decoder are linear layers, taking $\mathcal{O}(nd)$ time complexity and $\mathcal{O}(d)$ space complexity. Similarly, the location decoder with two modes with time complexity $\mathcal{O}(2nd)$ and space complexity $\mathcal{O}(2d)$. The GRU backbone has time and space complexities of $\mathcal{O}(nd^{2})$ and $\mathcal{O}(d^{2})$, respectively. For user VAE, we consider both input and hidden size to be $d$ for simplicity (hidden size is usually smaller than input size), and both time and space complexities are $\mathcal{O}(d^{2})$ because usually $d^{2}\gg dz$. In summary, MIRAGE has time complexity $\mathcal{O}(nd^{2})$ and space complexity $\mathcal{O}(d^{2})$. We train MIRAGE on a server with CPU: Intel(R) Xeon(R) Gold 6416H, GPU: NVIDIA GeForce RTX 4090). The training process takes 22.52s per epoch (converged at 50 epochs) and the generation takes 39.71s on the Istanbul dataset.
(a) Tokyo
(b) Istanbul
(c) NewYork
Figure 3. Return probability over different time distances
Appendix D Baselines
We present the details of the baseline methods below. Semi-Markov (Korolyuk et al., 1975) model utilizes exponential distribution with Gamma priors for time intervals and incorporates a Dirichlet prior to construct the transition matrix for Bayesian inference. TimeGeo (Jiang et al., 2016) designs a statistical EPR model integrated with temporal information including weekly home-based tour number, dwell rate, and burst rate to further characterize temporal choices on human mobility. RMTPP (Du et al., 2016) uses Recurrent Neural Networks to jointly model the time and mark dependency over history information. We consider the location ID as the mark here. ERTPP (Xiao et al., 2017) adopts distinct Recurrent Neural Networks to independently model the timing of the next event and its associated mark. THP (Zuo et al., 2020) incorporates the self-attention mechanism with the Hawkes Process to capture long-term dependencies in event sequence data. LSTM (Huang et al., 2015) learns the sequence patterns of human trajectory to predict both the location and time of the next event in a human trajectory. SeqGAN (Yu et al., 2017) introduces reinforcement learning into the GAN model to solve the sequence generation problem. MoveSim (Feng et al., 2020) is a GAN-based framework that integrates physical regularities and prior knowledge of human mobility in trajectory generation. LogNormMix (Shchur et al., 2019) defines an intensity-free neural TPP modeling the conditional probability density distribution of trajectory events as a log-normal mixture. We extend it as marked TPP to sample the time and location (represented as a mark) of the next event in a human trajectory. VOLUNTEER (Long et al., 2023) incorporates a two-layer VAE model with a temporal point process to capture the characteristics of human mobility from both group and individual views. DiffTraj (Zhu et al., 2024) utilizes a diffusion probabilistic model for the continuous location generation, it reconstructs and synthesizes geographic trajectories from white noise through a reverse trajectory denoising process. We choose the nearest POI to each generated GPS coordinate as the generated POI.
Appendix E Algorithms for the RecLoc Task
We present the details of the algorithms used to solve the location recommendation task below. BPR (Rendle et al., 2012) is a recommendation approach designed for implicit feedback. It operates by minimizing a pairwise ranking loss to learn user preferences on times effectively. DMF (Xue et al., 2017) incorporates a matrix factorization model with neural network architecture to the representations of users and locations. LightGCN (He et al., 2020) leverages neighborhood aggregation technique to learn user and location representations on a user-location interaction graph. MultiVAE (Liang et al., 2018) extends Variational Autoencoders (VAEs) to collaborative filtering for implicit feedback. NeuMF (He et al., 2017) proposes a neural network-based collaborative filtering technique and modeling user–location interaction function with non-linearities.
Appendix F Algorithms for the Nexloc Task
We present the details of the algorithms used to solve the next location recommendation task below. FPMC (Rendle et al., 2010) utilizes matrix factorization techniques to estimate the personalized transition matrix of POIs in user mobility trajectories. BERT4Rec (Sun et al., 2019) utilizes bidirectional self-attention network models human behavior sequences by employing a Cloze task approach. Caser (Tang and Wang, 2018) employs hierarchical and vertical CNNs to capture union-level sequential patterns and skip behaviors, enabling sequence-aware recommendation. SRGNN (Wu et al., 2019) designs a session-based recommendation model by utilizing the basic RNNs to predict the next location based on historical trajectories. SASRec (Kang and McAuley, 2018) employs a multi-head attention mechanism to make predictions based on the historical trajectory of users.
Appendix G Epidemic Simulation Settings
G.1. COVID-19 Simulation
We simulate the COVID-19 spreading with the SEIR model following recent works (Feng et al., 2020; Yuan et al., 2022; Lai et al., 2020). In particular, the simulation involves eight fundamental parameters: the close contact ratio ($c$), transmission period ($T$), incubation period ($T_{i}$), infection period ($T_{f}$), reproduction rate ($R_{0}$), transmission probability ($\beta$), infectious rate ($\alpha$), and recovery rate ($\gamma$). The detailed values of these parameters are shown in the Table 7.
Table 7. Parameters for COVID-19 simulation, where $d$ denotes days
Parameters $c$ $T$ $T_{i}$ $T_{f}$ $R_{0}$ $\beta$ $\alpha$ $\gamma$
Value 0.2 $5.8d$ $5.2d$ $11d$ 2.2 $R_{0}/T$ $1/T_{i}$ $1/T_{f}$
During the simulation, we assume that infected or exposed individuals contact with $s$ susceptible individuals connected by edges in the contact network each day. The probability of any two people with an edge in the contact network becoming close contact is $c$. The transmission probability $\beta$ is calculated by dividing the basic reproduction rate $R_{0}$ by the average duration (5.8 days) from onset to first medical visit and isolation. The infectious rate from the exposed state $\alpha$ is estimated as the reciprocal of the incubation period, which is the average time exposed but not infectious (5.2 days in (Lai et al., 2020)). The daily probability of transitioning to the removed state from infectious $\gamma$, is computed based on the average infection period (11 days in (Lai et al., 2020)). The formulas of the infection process are as follows:
G.2. Influenza Simulation
Similarly to the COVID-19 simulation, we simulate influenza spreading using the parameters suggested by (Brauer et al., 2019), where parameters transmission probability $\beta$ infectious rate $\alpha$ and removed rate $\gamma$ are shown in Table 8.
Table 8. Parameters for influenza simulation
Parameters $c$ $\beta$ $\alpha$ $\gamma$
Value 0.2 0.402 0.526 0.244
@@ G.2. Influenza Simulation
Appendix H Performance in Task-Based Evaluation on MSPE

### v2

@@ 5.2.2. Location Recommendation Task
(LocRec) suggests new (previously unvisited) locations for users by modeling user preferences on locations from trajectory data (Ye et al., 2010). Following the default setting of (Zhao et al., 2021), we transform a trajectory dataset into a set of (user, location, visit_count) triplets and then split them into training/valid/test datasets under a ratio of 8:1:1. To discount the impact of the specific techniques and metrics, we consider five popular recommendation algorithms, i.e., BPR (Rendle et al., 2012), DMF (Xue et al., 2017), LightGCN (He et al., 2020), MultiVAE (Liang et al., 2018), and NeuMF (He et al., 2017) (details in Appendix D), and report their performance on Mean Reciprocal Rank@N (MRR@N), Normalized Discounted Cumulative Gain@N (NDCG@N), hit@N (where N = 5 and 10). Finally, we compare the paired performance discrepancy between the real and generated trajectories (i.e., 30 paired results from five algorithms and six metrics each) using MAPE and MSPE.
@@ 5.2.3. Next Location Prediction Task
(NexLoc) forecasts a user’s next location in the future by learning the sequential mobility patterns from historical user trajectories (Noulas et al., 2012). Specifically, for one dataset, we chronologically split each trajectory into training/valid/test trajectories under a ratio of 8:1:1. We also consider five sequence prediction algorithms, i.e., FPMC (Rendle et al., 2010), BERT4Rec (Sun et al., 2019), Caser (Tang and Wang, 2018), SRGNN (Wu et al., 2019), and SASRec (Kang and McAuley, 2018) (details in Appendix E), and report their performance on MRR@N, NDCG@N, hit@N (where N = 5 and 10). Finally, we compare the paired performance discrepancy between the real and generated trajectories using MAPE and MSPE.
@@ 5.2.5. Epidemic Simulation Task
(EpiSim) simulates the epidemic spreading over a contact network characterizing the spatiotemporal contact patterns of user trajectories (Verma et al., 2021). The contact network is extracted from a trajectory dataset as a dynamic graph of users, and a directed edge from a user $p$ to a user $q$ represents that $p$’s visit precedes $q$’s visit to the same location during a day, indicating a potential chance of epidemic spreading from $p$ to $q$. Following the setting of recent works (Feng et al., 2020; Yuan et al., 2022; Lai et al., 2020), we adopt the Susceptible–Exposed–Infected–Recovered (SEIR) model. We consider the simulation of COVID-19 using the parameters suggested by (Feng et al., 2020; Yuan et al., 2022) and influenza using the parameters suggested by (Brauer et al., 2019) (details in Appendix F). We randomly select 50 individuals as exposed individuals, simulate the spread of the epidemic, and report the daily counts of Exposed, Infectious, and Recovered individuals as three metrics. We report the average results of 10 repeated simulations to discount the impact of the random selection of initially exposed individuals. Finally, we compare the performance discrepancy between the real and generated trajectories using MAPE and MSPE.
@@ 5.3. Statistical & Distributional Similarity
Table 2 shows the performance comparison of MIRAGE and baselines on the three datasets. We observe that MIRAGE consistently achieves the best performance with the lowest JSD, yielding 59.0%, 64.4%, and 67.7% improvement (reduction on JSD) over the best-performing baselines on TKY, IST, and NYC datasets, respectively. However, even though we selected four similarity metrics covering three different aspects of spatial (Distance and Radius), temporal (Interval), and user (DailyLoc) distributions, these metrics cannot fully reflect the ultimate utility of generated trajectories in supporting downstream tasks. For example, the consistent superiority of MIRAGE over baselines on these similarity metrics is still biased, because MIRAGE indeed underperforms some baselines in a few cases in our task-based evaluation, as we discuss below.
@@ 5.4. Task-Based Evaluation Performance
Table 3 shows the performance in our proposed task-based evaluation on MAPE (similar results on MSPE, shown in Appendix G). Each entry in this Table is the MAPE averaging over all metrics of all algorithms solving a task on a dataset, to average out the biases of individual techniques and metrics. For example, the MAPE of MIRAGE in the LocRec task on the TKY dataset is 0.3485, which is computed from 30 paired (real and generated) results from five algorithms and six metrics used in the LocRec task.
@@ 5.5. Ablation Study
We conduct an ablation study on our proposed method MIRAGE, considering the following four variants. MIRAGE-noTPP is a variant of MIRAGE without the neural TPP, where we define a regression task to predict the next inter-event from the time context vector $c_{i}^{\tau}$. MIRAGE-noEPR is a variant of MIRAGE without neural EPR, where we directly sample a location from a categorical distribution conditioned on the location context vector $c_{i}^{l}$. MIRAGE-noIMI is a variant of MIRAGE uses $t_{i}^{emb}$ and $k_{i}^{emb}$ (rather than $t_{i+1}^{emb}$ and $k_{i+1}^{emb}$) for sampling the next category and next POI. MIRAGE-TD is a variant of MIRAGE that uses temporal distances (rather than timestamps) in the exploration mode. Tables 4 and 5 show the results on similarities and task-based evaluation on MAPE, respectively (similar results on MSPE in Appendix G).
We observe that MIRAGE consistently outperforms MIRAGE-noTPP on both similarities and task-based evaluation by 63.8% and 30.0% (on average over tasks and datasets), respectively, showing the effectiveness of neural TPPs modeling the event stochasticity of human trajectories. Second, MIRAGE outperforms MIRAGE-noEPR in most cases, with an improvement of 50.3% and 17.8% on similarities and task-based evaluation, respectively, which verifies the usefulness of our neural EPR model. Third, MIRAGE outperforms MIRAGE-noIMI in general, by 24.8% and 18.1% on similarities and task-based evaluation, respectively, since the next location should depend on the next category and time, rather than the current ones. Finally, MIRAGE outperforms MIRAGE-TD in most cases, by 15.2% and 18.5% on similarities and task-based evaluation, respectively. Because timestamps offer more direct information for location sampling in the exploration mode compared to temporal distances. Our ablation studies systematically validate our key design choices.
In addition, to further show the utility of the EPR model, we plot the empirical returning probability of users over time, which is defined as the probability of a user returning to a location a certain period (temporal distance) after the user’s first presence at the location (Gonzalez et al., 2008; Yang et al., 2020a). Figure 2 shows the plots of both real and generated data on NYC. We see that the real trajectory exhibits strong periodicity, which can be well imitated by MIRAGE but not by MIRAGE-noEPR and baselines. The superiority of MIRAGE is also validated by the return probability over sequence lengths. Figure 3(a) shows the return probability of the whole trajectories of real and generated data. We selected top baselines SeqGAN, VOLUNTEER, and MoveSim for comparison. Figures 3(b), 3(c), and 3(d) show the detailed return probability w.r.t. the (percentage) length of trajectories of three datasets (100% corresponds to the whole trajectory). We see that our MIRAGE is more similar to the real data in terms of return trend across different (percentage) lengths. Additionally, Figure 4 shows the learnt return probability $p^{*}(l_{i+1}|\tau_{i+1},return)$ over different time distances (hour granularities in a week) from each dataset. We see a clear daily return pattern, implying that our model can effectively capture the periodicity encoded in human trajectories.
@@
(d) Semi-Markov
(e) Time Geo
(f) RMTPP
(g) ERTPP
(h) THP
(i) ActSTD
(j) LogNormMix
(k) LSTM
(l) SeqGAN
(m) MoveSim
(n) VOLUNTEER
(o) DiffTraj
@@
(a) Return Probability Distribution on Three Datasets
(b) Return Probability Distribution over Sequence on Tokyo
(c) Return Probability Distribution over Sequence on Istanbul
(d) Return Probability Distribution over Sequence on NewYork
Figure 3. Returning Probability over Sequence Length
(a) Tokyo
(b) Istanbul
(c) NewYork
Figure 4. Return probability over different time distances
@@ 6. Conclusions and Future Works
In this paper, by revisiting the existing human trajectory generative models, we identify their limitations in focusing on the summary statistics and distributional similarities between real and generated trajectories, which could lead to intrinsic biases in both generative model design and benchmarks of the generated trajectories. Against this background, we propose MIRAGE, a huMan-Imitative tRAjectory GenErative model designed as an intensity-free neural Temporal Point Process integrating a neural Exploration and Preferential Return model to imitate the human decision-making process in trajectory generation. Meanwhile, we also propose a comprehensive task-based evaluation protocol to systematically benchmark trajectory generative models on four typical downstream tasks, integrating multiple techniques and evaluation metrics for each task, to assess the ultimate utility of the generated trajectories. The evaluation results show that MIRAGE-generated trajectory data not only achieves the best statistical and distributional similarities with 59.0-67.7% improvement but also yields the best performance in the task-based evaluation with 10.9-33.4% improvement.
@@ Acknowledgements.
This project has received funding from the University of Macau (MYRG2022-00048-IOTSC), the Science and Technology Development Fund, Macau SAR (0047/2022/A1, 001/2024/SKL), and Jiangyin Hi-tech Industrial Development Zone under the Taihu Innovation Scheme (EF2025-00003-SKL-IOTSC). This work was performed in part at SICC which is supported by SKL-IOTSC, University of Macau.
@@ Appendix A Biases in downstream tasks
We show the performance variation of different methods solving the same downstream task in our task-based evaluation on MAPE in Table 6. Taking MoveSim and VOLUNTEER as examples, MoveSim is better than VOLUNTEER when benchmarking using FPMC or SASRec in this task, while we have the opposite results if benchmarking using BERT4Rec, Caser, or SRGNN.
@@
Appendix B The Complexity of MIRAGE
@@
Let $n$ denote the sequence length, $d$ is the GRU hidden state size, and $z$ is the latent size of user VAE. Our time decoder with intensity-free TPP and category decoder are linear layers, taking $\mathcal{O}(nd)$ time complexity and $\mathcal{O}(d)$ space complexity. Similarly, the location decoder with two modes with time complexity $\mathcal{O}(2nd)$ and space complexity $\mathcal{O}(2d)$. The GRU backbone has time and space complexities of $\mathcal{O}(nd^{2})$ and $\mathcal{O}(d^{2})$, respectively. For the user VAE, we consider both input and hidden size to be $d$ for simplicity, and both time and space complexities are $\mathcal{O}(d^{2})$ as usually $d^{2}\gg dz$. In summary, MIRAGE has time complexity $\mathcal{O}(nd^{2})$ and space complexity $\mathcal{O}(d^{2})$.
@@
Appendix C Baselines
Semi-Markov (Korolyuk et al., 1975) model uses exponential distribution with Gamma priors for time intervals and incorporates a Dirichlet prior to construct the transition matrix for Bayesian inference. TimeGeo (Jiang et al., 2016) designs a statistical EPR model integrated with temporal information including weekly home-based tour number, dwell rate, and burst rate to further characterize temporal choices on human mobility. RMTPP (Du et al., 2016) uses Recurrent Neural Networks to jointly model the time and mark (location ID) dependency over history information. ERTPP (Xiao et al., 2017) adopts distinct RNNs to independently model the timing of the next event and its associated mark. THP (Zuo et al., 2020) incorporates the self-attention mechanism with the Hawkes Process to capture long-term dependencies in event sequence data. LSTM (Huang et al., 2015) learns the sequence patterns of human trajectory to predict both the location and time of the next event in a human trajectory. SeqGAN (Yu et al., 2017) introduces reinforcement learning into the GAN model to solve the sequence generation problem. MoveSim (Feng et al., 2020) is a GAN-based framework that integrates physical regularities and prior knowledge of human mobility in trajectory generation. ActSTD (Yuan et al., 2022) enhances the dynamic modeling of individual trajectories by utilizing neural ordinary equations. LogNormMix (Shchur et al., 2019) defines an intensity-free neural TPP modeling the conditional probability density distribution of trajectory events as a log-normal mixture. We extend it as marked TPP by sampling the time and location (i.e., mark) of the next event in a trajectory. VOLUNTEER (Long et al., 2023) incorporates a two-layer VAE model with a temporal point process to capture the characteristics of human mobility from both group and individual views. DiffTraj (Zhu et al., 2024) uses a diffusion probabilistic model for continuous location generation. We choose the nearest POI to each generated GPS coordinate as the generated POI.
Appendix D Algorithms for the RecLoc Task
BPR (Rendle et al., 2012) is a recommendation approach designed for implicit feedback. It operates by minimizing a pairwise ranking loss to learn user preferences on times effectively. DMF (Xue et al., 2017) incorporates a matrix factorization model with neural network architecture to the representations of users and locations. LightGCN (He et al., 2020) leverages neighborhood aggregation technique to learn user and location representations on a user-location interaction graph. MultiVAE (Liang et al., 2018) extends Variational Autoencoders (VAEs) to collaborative filtering for implicit feedback. NeuMF (He et al., 2017) proposes a neural network-based collaborative filtering technique and modeling user–location interaction function with non-linearities.
Appendix E Algorithms for the Nexloc Task
FPMC (Rendle et al., 2010) uses matrix factorization techniques to estimate the personalized transition matrix of POIs in user mobility trajectories. BERT4Rec (Sun et al., 2019) uses bidirectional self-attention network models human behavior sequences by employing a Cloze task approach. Caser (Tang and Wang, 2018) employs hierarchical and vertical CNNs to capture union-level sequential patterns and skip behaviors, enabling sequence-aware recommendation. SRGNN (Wu et al., 2019) designs a session-based recommendation model by utilizing the basic RNNs to predict the next location based on historical trajectories. SASRec (Kang and McAuley, 2018) employs a multi-head attention mechanism to make predictions based on the historical trajectory of users.
Appendix F Epidemic Simulation Settings
We simulate the COVID-19 spreading with the SEIR model following recent works (Feng et al., 2020; Yuan et al., 2022; Lai et al., 2020). In particular, the simulation involves eight fundamental parameters: the close contact ratio ($c$), transmission period ($T$), incubation period ($T_{i}$), infection period ($T_{f}$), reproduction rate ($R_{0}$), transmission probability ($\beta$), infectious rate ($\alpha$), and recovery rate ($\gamma$). Table 9 shows the parameter values.
Table 9. Parameters for COVID-19 simulation ($d$: day)
Parameters $c$ $T$ $T_{i}$ $T_{f}$ $R_{0}$ $\beta$ $\alpha$ $\gamma$
Value 0.2 $5.8d$ $5.2d$ $11d$ 2.2 $R_{0}/T$ $1/T_{i}$ $1/T_{f}$
During the simulation, we assume that infected or exposed individuals contact with $s$ susceptible individuals connected by edges in the contact network each day. The probability of two people with an edge in the contact network becoming close contact is $c$. The transmission probability $\beta$ is calculated by dividing the basic reproduction rate $R_{0}$ by the average duration (5.8 days) from onset to first medical visit and isolation. The infectious rate from the exposed state $\alpha$ is estimated as the reciprocal of the incubation period, which is the average time exposed but not infectious (5.2 days in (Lai et al., 2020)). The daily probability of transitioning to the removed state from infectious $\gamma$, is computed based on the average infection period (11 days in (Lai et al., 2020)). The the infection process are as follows:
Similarly to the COVID-19 simulation, we simulate influenza spreading using the parameters suggested by (Brauer et al., 2019), where the transmission probability $\beta$ infectious rate $\alpha$ and removed rate $\gamma$ are shown in Table 10.
Table 10. Parameters for influenza simulation
Parameters $c$ $\beta$ $\alpha$ $\gamma$
Value 0.2 0.402 0.526 0.244
Appendix G Performance on MSPE